What is the correct process of solving:
9x + 10 = 29 - 13x ?

Answer by queenofit(120)

Ok, first let me tell you that I like to think of the numbers and the variables being on seperate "teams". You have to be fair, so what you "do" to one team, you must do to the other team. The teams want to be on opposite sides of the = sign, but remember when they cross the = sign, the neg changes to pos and the pos changes to neg. I will work it out long so you will see exactly what is going on. so lets get to work.
9x+10=19-13x lets do the -13x first.
9x+13x+10=19-13x+13x I did this to move the -13x to the correct side.
9x+13x+10=19 the -13 is gone from the right side now.
9x+13x+10-10=19-10 now we do the same with the 10
9x+13x=19-10 see how we moved the teams to the correct side?
22x=9 combine the like terms
but you are not solving for 22x, you want x, so divide each side by 22
22x/22=9/22
x=9/22
do you understand? you can cut this down by just moving the numbers and variables across the = sign and changing them from neg to pos or pos to neg depending on the sign they had to begin with. when you move them across, change them to the opposite sign.
x=9/22 may look strange to you and your instructor may want it in a different form such as a decimal, or have you reduce if you can...
I hope this helped, let me know if you need anything further.